OpenStudy (anonymous):

i have (x^2+6x)+(y^2+4y)+(z^2-2z)=1 how do i complete the square

OpenStudy (campbell_st):

if you have a quadratic \[x^2 + bx + ? \] to complete the square you need to find \[\frac{b}{2} \] then square it.... and add it to both sides of the equation so \[x^2 + bx + (\frac{b}{2})^2 = (x + \frac{b}{2})^2\] hope this helps

OpenStudy (jdoe0001):

freddy_eighty7 do you know what a \(\bf \text{perfect square trinomial}\) is ?

OpenStudy (anonymous):

so i do each one at a time? (x,y,z) i don't know about the perfect square

OpenStudy (anonymous):

\[ x^2+bx = \left(x+\frac b 2\right)^2-\left(\frac b 2\right)^2 \]

OpenStudy (campbell_st):

yes... some process for each

OpenStudy (jdoe0001):

hmmmm.... well... you'd need to, otherwise completing the square won't be of much use http://www.youtube.com/watch?v=7rffjYDxW5I

OpenStudy (campbell_st):

but make sure you add the term (b/2)^2 to both sides of the equation... to keep it in balance

OpenStudy (anonymous):

so the pattern I'm seeing is half the x, then square it, and that becomes the b to where you also add to the other side of the = to even it out, yes?

OpenStudy (anonymous):

and its always +?

OpenStudy (campbell_st):

well if you have a perfect square then milled term is always double to product of the 2 terms in the binomial... the last term is always positive... \[(m - n)2 = m^2 - 2mn + n^2\] the middle term always has the same sign as the binomial...

OpenStudy (anonymous):

i think i got it, thank you ill prob bug you more later since your answering the way i want to by giving me the base formula